The AASTEX Package for Manuscript Preparation

Contents

1 Introduction 1

2 AASTEX article markup 1

2.1 Preamble . . . . . . . . . . . . . . . . . . 1 2.1.1 Getting started . . . . . . . . . . . 2 2.1.2 Defining new commands . . . . . . 2 2.1.3 Editorial information . . . . . . . . 2 2.1.4 Short comment on title page … 2 2.1.5 Running heads . . . . . . . . . . . 2

2.2 Starting the main body . . . . . . . . . . 3 2.3 Title and author information . . . . . . . 3 2.4 Abstract . . . . . . . . . . . . . . . . . . . 3 2.5 Keywords . . . . . . . . . . . . . . . . . . 3 2.6 Comments to editors . . . . . . . . . . . . 3 2.7 Sections . . . . . . . . . . . . . . . . . . . 4 2.8 Figure and table placement . . . . . . . . 4 2.9 Acknowledgments . . . . . . . . . . . . . . 4 2.10 Appendices . . . . . . . . . . . . . . . . . 4 2.11 Equations . . . . . . . . . . . . . . . . . . 4 2.12 Citations and bibliography . . . . . . . . . 5

2.12.1 The thebibliography environment 5 2.12.2 Specifying bibliographic and cita-

tion information . . . . . . . . . . 5 2.12.3 The references environment … 5 2.12.4 Abbreviations for journal names . 6

2.13 Figures . . . . . . . . . . . . . . . . . . . 6 2.13.1 Electronic art . . . . . . . . . . . . 6 2.13.2 Figure captions . . . . . . . . . . . 7

2.14 Tables . . . . . . . . . . . . . . . . . . . . 7 2.14.1 The deluxetable environment . . 7 2.14.2 Preamble to the deluxetable … 7 2.14.3 Content of the deluxetable … 8 2.14.4 The table environment . . . . . . 8 2.14.5 Table footnotes . . . . . . . . . . . 9

2.15 Miscellaneous . . . . . . . . . . . . . . . . 9 2.15.1 Ionic species and chemical bonds . 9 2.15.2 Fractions . . . . . . . . . . . . . . 9 2.15.3 Celestial objects . . . . . . . . . . 10 2.15.4 Hypertext constructs . . . . . . . . 10

2.16 Concluding the file . . . . . . . . . . . . . 10

3 Style options 10 3.1 Preprint format . . . . . . . . . . . . . . . 10

3.2 Two-column format . . . . . . . . . . . . 11

4 Additional documentation 11

5 Acknowledgments 11

1. Introduction

The American Astronomical Society (AAS) has developed an author markup package called AASTEX that assists authors in preparing manuscripts that are intended for submission to the AAS-affiliated journals. AASTEX can be used for other journals as well, some of which accept AASTEX manuscripts.

The most important aspect of the AASTEX package is that it defines the set of commands (called markup) that can be used to identify the structural elements of a paper. When articles are marked up using this set of standard commands, the papers may then be submitted electronically to the editorial offices and fed into electronic production of the journals.

This guide contains basic instructions for creating manuscripts using the AASTEX markup package. Authors are expected to be familiar with the editorial requirements of the journals so that they can make appropriate submissions; they should also have at least a rudimentary knowledge of L^{A}TEX (for instance, knowing how to set up equations using L^{A}TEX commands). A number of useful publications about L^{A}TEX (and TEX) are listed in the reference section of this guide.

Authors who wish to submit papers electronically to the ApJ, AJ, or PASP are strongly encouraged to use the AASTEX markup package as described in this guide.

2. AASTEX article markup

This section describes the commands in the AASTEX package that an author might enter in a manuscript that is being prepared for electronic submission to one of the journals. The commands are described in roughly the same order as they would appear in a manuscript. The reader will also find it helpful to examine the sample file (sample.tex) that is distributed with the package. Authors are reminded to check the electronic submission procedures and the instructions to authors for the journals to which they submit their papers.

2.1. Preamble

In L^{A}TEX manuscripts, the preamble is that portion of the file before the \begin{document} command.

2.1.1. Getting started

The first piece of markup in the manuscript declares the overall style of the document. Any commands that appear before this markup will be ignored.

     \documentclass{aastex}

The above example specifies the document class as aastex, with the default style, manuscript, in effect. The paper copy produced by this style file will be doublespaced. Any tables included in the main body of the manuscript will also be double-spaced.

Other styles are available; they are discussed in Section 3.

2.1.2. Defining new commands

AASTEX allows authors to define their own commands with L^{A}TEX’s \newcommand. (Do not use the plain TEX \def command.) Authors’ \newcommand definitions must be placed in the document preamble.

In general, author-defined commands that are abbreviations or shorthands are acceptable and can be translated, e.g.,

     \newcommand{\lte}{local thermodynamic
     equilibrium}

However, abbreviations that attempt to define new symbols by using L^{A}TEX commands for repositioning text or symbols tend to cause problems during the production of the article.

Specifically, author-defined commands that use any of the commands listed below are apt to be problematic in the conversion procedures after the manuscript is submitted.

     \hskip, \vskip, \raise, \raisebox, \lower,
     \rlap, \kern, \lineskip, \char, \mathchar,
     \mathcode, \buildref, \mathrel, \baselineskip

Consequently, authors are strongly discouraged from employing them.

Extra symbols are defined for AASTEX, some specifically for an astronomical context, and others more broadly used in math and physics. In particular, the AMS has additional symbol fonts that are available in a standard L^{A}TEX package (amssymb). All of these symbols are depicted in the additional tables supplied with this manual.

Authors are advised to consult these tables to see whether a symbol already exists. If it does, you should just use the corresponding markup command; these commands are reserved. You should not redefine an existing command name. When one of these commands is encountered in an electronic manuscript, authors’ redefinitions are ignored when the manuscript is translated.

2.1.3. Editorial information

A number of markup commands are available for the editorial offices to record the publication history and slug-line data for each manuscript. These commands are for use in the editorial and production offices only.

                                                                 \received{receipt date}

\revised{revision date} \accepted{acceptance date}

                                                                 \ccc{code}
                                                                 \cpright{type}{year}

Copyright information is specified through the commands \cpright and \ccc. The “type” of copyright and the corresponding year are given in \cpright; valid copyright types are as follows.

AAS Copyright has been assigned to the AAS
ASP Copyright has been assigned to the ASP
PD  The article is in the public domain
none No copyright is claimed for the article

The copyright type is case sensitive, so the type string must be entered exactly as given above. The Copyright Clearing Center code may be given in the \ccc command; the code is taken as regular text, so any special characters, notably ‘$’, must be properly specified.

2.1.4. Short comment on title page

Authors who wish to include a short remark on the title page, such as the name and date of the journal in which an article has been scheduled, may do so with the following command.

                                                                 \slugcomment{text}

In the manuscript style, such comments appear on the title page after the title and authors; in the preprint style, they are placed at the upper right corner of the title page.

2.1.5. Running heads

Authors are invited to supply running head information. There are generally two different kinds of data in running heads; the left head contains an author list (last names, possibly truncated as “et al.”), while the right head is an abbreviated form of the paper title. This running head information will not appear on the L^{A}TEXprinted page but will be transmitted to the copy editor for inclusion in the published version.

                                                                 \shorttitle{text}
                                                                 \shortauthors{text}

Editors and publishers impose varying requirements on the brevity of these data. A good rule of thumb is to limit the list of authors to three, or use “et al.”, and to limit the short form of the title to 40–45 characters. The editors may choose to modify the author-supplied running heads.

2.2. Starting the main body

None of the markup that appears in the preamble actually typesets anything; it is only a control section. The author must include a

     \begin{document}

command to identify the beginning of the main textual portion of the manuscript.

2.3. Title and author information

Use the \title and \author commands to specify title and author information. Specify the author’s principle affiliation with the command \affil. Each \author command should be followed by a corresponding \affil and optional \email command.

     \title{text}
     \author{name(s)}

\affil{affiliation}

     \affil{address}

\email{email address}

     \and

Line breaks are permitted in the title if the author wishes to specify them with the \\ command. Long titles will be broken automatically, so the \\ markup is not required. If the title is explicitly broken over several lines, the preferred style for titles in AAS and ASP journals is the so-called “inverted pyramid” style. In this style, the longest line is the first (or top) line, and each succeeding line is shorter. The text of the title should be entered in mixed case; it will be published in upper case or mixed case according to the style of the publication to which the paper is submitted. Footnotes are permissible in titles; be careful to ensure that alternate affiliations (see below) are properly numbered if a footnote to the title is specified.

Authors’ names are given in \author commands, and should be entered in mixed case. Names that appear together in the author list for authors who have the same primary affiliation should be specified in a single \author command. Each author group (\author command) should be followed by an \affil command, giving the principle affiliation of that author. Physical and postal address information for the specified institution may be included with the \affil. The address can be broken over several lines using the \\ command to indicate the line breaks. Usually, however, postal information will fit on one line. When there is more than one \author command, the last one should be preceded by the \and command.

When there is a lengthy author list, all authors’ names may be specified in a single \author command, with affiliations specified using the \altaffilmark mechanism described below. In this case, no \affil commands would be used, and in print the affiliations would all be listed in a footnote block at the bottom of the title page. Authors often have affiliations in addition to their principle employer. These alternate affiliations are spec-

                                                      ified with the \altaffilmark and \altaffiltext com-

mands. These behave like the \footnotemark and \footnotetext commands of L^{A}TEX, except there are no optional arguments in AASTEX. \altaffilmark is appended to authors’ names in the \author list and generates superscript identification numbers. The text for the individual alternate affiliations is generated by the

                                                      \altaffiltext command.
                                                                 \altaffilmark{key number(s)}
                                                                 \altaffiltext{numerical key}{text}

It is up to the author to make sure that each key number in his or her \altaffilmark matches the numerical key for the corresponding \altaffiltext.

2.4. Abstract

The paper abstract should be enclosed in the abstract environment.

                                                                 \begin{abstract}

abstract text

                                                                 \end{abstract}

2.5. Keywords

Keywords, subject headings, etc., are accommodated as a single piece of text.

                                                                 \keywords{text}

If authors supply keywords, they must be delimited by whatever punctuation is required by the journal. They should be specified in alphabetical order. The \keywords command supplies the proper leading text (“Keywords:”, “Subject headings:”, etc.), according to journal style.

2.6. Comments to editors

Authors may make notes or comments to the copy editor with the \notetoeditor command.

                                                                 \notetoeditor{text}

This command behaves like a footnote and appears on the bottom of the page. Output to the printed page is produced only in the manuscript style.

2.7. Sections

The manuscript style for AASTEX manuscripts supports four levels of section headings.

     \section{heading}
     \subsection{heading}
     \subsubsection{heading}
     \paragraph{heading}

Section headings should be given in upper case or mixed case depending on the style of the journal. Note that these commands delimit sections by marking the beginning of each section; there are not separate commands to identify the ends.

2.8. Figure and table placement

Generally figures and tables are not “placed” in the text of the document where an author would like them to appear physically. However, authors may indicate to the editors the preferred placement of these items using the \place* commands.

     \placetable{key}
     \placefigure{key}

The \place*{key} commands are similar to the \ref command in LaTEX and require corresponding \label commands to link them to the proper elements.

When used in the manuscript style, the \place* commands will print a short message to the editor about figure and table placements. In the other styles, nothing is printed.

2.9. Acknowledgments

AASTEX manuscript styles support an acknowledgments section.

     \acknowledgments

In the AASTEX styles, acknowledgments are set off from the concluding main body text by extra vertical space, with no heading or type size change.

2.10. Appendices

When one or more appendices are needed in a paper, the point where the main body text ends and the appendix begins must be marked with the \appendix command.

\section{body section}

     \appendix

\section{appendix section}

The \appendix command takes care of a number of internal housekeeping concerns, such as identifying sections with letters instead of numerals, resetting the equation counter, etc. Note that the \appendix command takes no arguments. Sections in the appendix must be headed with \section commands, as described earlier.

2.11. Equations

Display equations can be typeset in L^{A}TEX in many ways; the following three are probably of greatest use in AASTeX.

                                                                 \begin{displaymath}
                                                                 \end{displaymath}
                                                                 \begin{equation}
                                                                 \end{equation}
                                                                 \begin{eqnarray}
                                                                 \end{eqnarray}

The displaymath environment will break out a single, unnumbered formula. The equation environment does the same thing, except that the equation is autonumbered by L^{A}TEX. To set several formulæ in which vertical alignment is required or to display a long equation across multiple lines, use the eqnarray environment. Each line of the eqnarray will be numbered unless a \nonumber command is inserted before the equation delimiter (\\). L^{A}TEX’s equation counter is not incremented when \nonumber is used.

Authors may occasionally wish to group related equations together and identify them with letters appended to the same equation number. When this is desired, such related equations should still be set in equation or eqnarray environments, whichever is appropriate, and the grouping should then be placed within the mathletters environment.

                                                                 \begin{mathletters}

equation or eqnarray

                                                                 \end{mathletters}

It is possible to override L^{A}TEX’s automatic numbering within the equation or eqnarray environments using

                                                                 \eqnum{text}

When \eqnum is specified inside an equation environment, or on a particular equation within an eqnarray, the text supplied as an argument to \eqnum is used as the equation identifier. L^{A}TEX’s equation counter is not incremented when \eqnum is used. \eqnum must be used inside the environment.

If, as a consequence of the use of \eqnum or \nonumber, L^{A}TEX’s equation counter gets out of synch with the author’s intended sequence, the counter may be reset to a particular value.

                                                                 \setcounter{equation}{number}

The equation counter should be set to the number corresponding to the last equation that was formatted; therefore, it is most appropriate for this command to appear immediately after an equation or eqnarray environment ends. The command must be used outside the math environments.

The eqsecnum style file can also be used to modify the way equations are numbered. See Section 3 for details.

2.12. Citations and bibliography

Two options are available for symbolically marking citations and formatting reference lists: the standard L^{A}TEX thebibliography environment, and the AASTEX references environment. Authors are strongly encouraged to use thebibliography in electronic submissions.

Please note that the bibliographic data supplied by the author in the reference list must conform to the standards of the journal. We have elected not to burden authors with tedious markup commands to delimit the bibliographic fields — many of the journals that accept AASTEX agreed to reduce typographic overhead (bold, italic, etc.) in reference lists (Abt 1990). It is the responsibility of the author to arrange these fields in the proper order with the correct punctuation; the information will be typeset in roman with no size or style changes.

2.12.1. The thebibliography environment

The preferred method for reference management is to use L^{A}TEX’s thebibliography environment, marking citations in the body of the paper with \citep or \citet and associating references with them via \bibitem. The \cite-\bibitem mechanism associates citations and references symbolically while maintaining proper citation syntax within the paper. In the \bibitem command, the author specifies citation data inside square brackets and a citation key in curly braces for each reference.

     \begin{thebibliography}{dummy}

\bibitem[cite data]{key} bibliographic data

        .
        .
     \end{thebibliography}

Note that the argument dummy to the environment is not used in the AASTEX package, but is consistent with the syntax of standard L^{A}TEX. It is acceptable to simply insert an empty pair of curly braces after the

\begin{thebibliography} command. The \bibitem

command is described in detail in the next section.

2.12.2. Specifying bibliographic and citation information

The scheme used in AASTEX for marking citations and for specifying bibliographic data allow us to utilize the features of the natbib package (Daly 1998). The natbib package re-implements L^{A}TEX’s \cite command, and offers greater flexibility for managing citations in the author-year form.

All of the bibliographic data are defined in \bibitem commands.

\bibitem[author(year)]{key} bibliographic data

The square-bracketed argument of \bibitem contains the author portion of the citation, followed by the year set off in parentheses. The parentheses are important — please don’t forget them. The curly-bracketed argument key is the code name by which the reference is known.

When placing citations in the text, the author should use either a \citep or a \citet command.

                                                                 \citep[extra]{key(s)}\\

\citet[extra]{key(s)}

The \citep command produces a citation that is entirely set off by parentheses, e.g. (Cox 1995), while \citet permits the author’s name to form part of your text, e.g. Cox (1995). The classic L^{A}TEX \cite command behaves like

                                                      \citet.

The citation key corresponds to a key in a \bibitem command. During processing, information from the cite data is inserted in the text at the location of the \cite command. Multiple citation keys are separated

                                                      by commas, e.g., \citep{knuth84,cox95,lamport94}.

\citep and \citet each take optional arguments that specify extra text that is appended to the citation label, per standard L^{A}TEX behavior.

AASTEX uses a full natbib implementation. There are more syntax options than we have described here, and these support the whole range of arcane possibilities for composing citations; the price is complexity. Authors who need more flexibility should become familiar with the natbib documentation. The syntax discussed above should be sufficient for the vast majority of cases.

It is not possible to use \bibitem within AASTEX’s references environment (section 2.12.3), nor will \cite commands work properly in the main body if \bibitem commands are absent.

2.12.3. The references environment

Some authors might prefer to enter citations directly into the body of an article. If so, the references environment may be used to format the reference list. The references environment simply sets off the list of references and adjusts spacing parameters.

                                                                 \begin{references}

\reference{key}bibliographic data

                                                                    .
                                                                    .
                                                                 \end{references}

While the references environment remains supported in v5.0, we anticipate that authors will prefer the stronger capabilities of the standard L^{A}TEX

                                                      thebibliography commands as extended by natbib.

2.12.4. Abbreviations for journal names

There are markup commands for many of the most frequently-referenced journals so that authors may use the markup rather than having to look up a particular journal’s abbreviation. In principle, all the journals should be using the same abbreviations, but it is fair to anticipate some changes in the specific abbreviations before a system is finally settled on. As long as these commands are kept up to date, authors need not be concerned about such editorial preferences and changes. A listing of the current abbreviations appears in Table 1.

2.13. Figures

2.13.1. Electronic art

For manuscripts that are submitted electronically, the electronic figures themselves are generally sent as individual files to the editors and are not included as part of the main text, although the legends or captions for these files are included using the markup described in Section 2.13.2. Detailed information on preparing and submitting electronic art is available in the submissions instructions for the various journals.

At this time, the most widely used means of including figures or other types of nontextual data in electronic manuscripts is to generate such files as Encapsulated PostScript files.^{1}

If an author wishes PostScript figures to be included in the pages being produced, it is necessary that the graphics files being included conform to the Encapsulated PostScript standard (Adobe 1999). The author must also have an appropriate DVI translator, one that targets PostScript output devices.

There are several commands to use when including EPS files in AASTEX manuscripts. They should be placed within the figure environment.

\begin{figure}
\figurenum{text}
\epsscale{num}
\plotone{epsfile}
\plottwo{epsfile}{epsfile}
\caption{text}
\end{figure}

When \figurenum is specified inside a figure environment, the text supplied as an argument to \figurenum is used as the figure identifier. L^{A}TEX’s figure counter is not incremented when \figurenum is used. \figurenum must be used inside the figure environment.

\plotone inserts the graphic in the named EPS file, scaled (in both dimensions) so that the horizontal dimension fits in the body text width; the vertical dimension Table 1: Abbreviations for Journal Names

\araa  Annual Review of Astronomy

and Astrophysics

\ao  Applied Optics
\aj  Astronomical Journal
\azh  Astronomicheskii Zhurnal
\aap  Astronomy and Astrophysics
\aapr  Astronomy and Astrophysics Reviews
\apj  Astrophysical Journal
\apjl  , Letters to the Editor
\apjs  , Supplement Series
\aplett  Astrophysics Letters
\apspr  Astrophysics Space Physics Research
\apss  Astrophysics and Space Science
\aaps  , Supplement Series
\baas  Bulletin of the AAS
\bain  Bulletin Astronomical Inst. Netherlands
\gca  Geochimica Cosmochimica Acta
\grl  Geophysics Research Letters
\iaucirc IAU Circular
\jcp  Journal of Chemical Physics
\jgr  Journal of Geophysics Research
\jrasc  Journal of the RAS of Canada
\memras  Memoirs of the RAS
\mnras  Monthly Notices of the RAS
\nat  Nature
\nphysa  Nuclear Physics A
\physscr Physica Scripta
\pra  Physical Review A
\prb  Physical Review B
\prc  Physical Review C
\prd  Physical Review D
\pre  Physical Review E
\prl  Physical Review Letters
\physrep Physics Reports
\planss  Planetary Space Science
\procspie Proceedings of the SPIE
\pasj  Publications of the ASJ
\pasp  Publications of the ASP
\qjras  Quarterly Journal of the RAS
\skytel  Sky and Telescope
\solphys Solar Physics

1PostScript is a registered trademark of Adobe Systems Incorporated.

is scaled to maintain the aspect ratio. \plottwo inserts two plots next to each other. Scale factors are determined automatically from information in the EPS file.

The automatic scaling may be overridden with the command \epsscale{num}, where num is in decimal units, i.e., 0.80.

2.13.2. Figure captions

Figure captions, or legends, should be included as a group at the end of the manuscript using either the \figcaption command or a figure environment for each figure in the paper, whether or not electronic art is included. For electronic submissions, use of \figcaption is preferred.

     \figcaption[filename]{text\label{key}}

The optional argument, filename, can be used to identify the file for the corresponding figure; text refers to the caption for that figure. The author may provide a \label with a unique key for cross-referencing purposes.

When the \figcaption command is used, the figure identification, e.g., “Figure 1,” is generated automatically by the command itself, so there is no need to key this information. There is an upper limit of seven figure captions per page. Footnotes are not supported for figures.

2.14. Tables

There is support in the AASTEX package for tables via two mechanisms: L^{A}TEX’s standard table environment and the deluxetable environment, which facilitates the formatting of lengthy tabular material. Tables may be marked up using either mechanism, although use of deluxetable is preferred. Authors should not use the L^{A}TEX tabbing environment or the \hline command when preparing electronic submissions.

L^{A}TEX permits the preparation of fairly complex tables with arbitrary spacing, straddle heads, rules, and the like. Authors who need to specify complicated column headings and so forth are advised to consult the L^{A}TEX manual (Lamport 1994). Most of the capabilities outlined there are applicable to the AASTEX table preparation environments.

2.14.1. The deluxetable environment

Authors are encouraged to use the deluxetable environment to format their tables since it automatically handles many formatting tasks, including table numbering and insertion of horizontal rules. It also provides mechanisms for breaking tables and controlling width and vertical spacing that are unavailable in the L^{A}TEX tabular environment.

The deluxetable environment is delimited by L^{A}TEX’s familiar \begin and \end constructs. The content consists of preamble commands and table data, the latter

                                                      delimited by \startdata and \enddata.
                                                                 \begin{deluxetable}{cols}

preamble commands

                                                                 \startdata

table data

                                                                 \enddata
                                                                 \end{deluxetable}

where cols specifies the justification for each column. One of the letters ‘l’, ‘c’, ‘r’, or ‘p’ is given for each column, indicating flush left, centered, flush right, or justified text. Authors are referred to the L^{A}TEX manual (Lamport 1994) for further information.

2.14.2. Preamble to the deluxetable

There are several items in the deluxetable environment that must be given before the data for the table; they constitute the preamble to the environment.

                                                                 \tabletypesize{font size command}
                                                                 \rotate
                                                                 \tablewidth{dimen}
                                                                 \tablenum{text}
                                                                 \tablecolumns{num}
                                                                 \tableheadfrac{num}
                                                                 \tablecaption{text\label{key}}
                                                                 \tablehead{text}

If a table is too wide for the printed page, it is permissible to change the font size within the deluxetable environment using the \tabletypesize command with the font size change commands \small (11pt), \footnotesize (10pt), or \scriptsize (8pt).

You can force a table to be set in landscape orientation using the \rotate command. The default is for AASTEX itself to determine if landscape orientation is required. The width of a deluxetable is defined by \tablewidth; the default width, if this command is omitted, is the width of the body text. The table can be set to its natural width by specifying a dimen of 0pt. Long tables may have a natural width that is different for each page. The natural width for each page will be printed to the log file during processing; authors may then define a fixed table width based on this information, giving the tables a more uniform appearance across the pages.

It is possible to override L^{A}TEX’s automatic numbering within the deluxetable environment. When \tablenum is specified inside a deluxetable preamble, the text supplied as an argument to \tablenum is used as the table identifier. L^{A}TEX’s equation counter is not incremented when \tablenum is used.

The caption (actually, the title) of the table is specified

                                                      in \tablecaption. The text of \tablecaption should

be brief; explanatory notes should be specified in the end notes to the table (see \tablecomments below). If the caption does not appear centered above the table after processing, then specify the width of the table explicitly in the \tablewidth command and rerun L^{A}TEX on the file. The author may provide a \label in the caption with a unique key for cross-referencing purposes.

Column headings are specified with \tablehead. Within \tablehead, each column heading can be given in a \colhead, which will ensure that the heading is centered on the natural width of the column; this is the typical disposition of column headings, and the use of \colhead is encouraged. There should be a heading for each column, so that there are as many \colhead commands in the \tablehead as there are data columns.

     \tablehead{
     \colhead{heading} & \colhead{heading}}

If more complicated column headings are required, any valid tabular command that constitutes a proper head line in a L^{A}TEX table may be used. Consult the L^{A}TEX manual (Lamport 1994) for details.

The \tablecolumns{num} command is necessary if the author has multi-line column headings produced by \tablehead or other L^{A}TEX commands and is using either the \cutinhead or \sidehead markup (see below). The num is set to the true number of columns in the table. The command must come before the \startdata command.

It is possible that a complicated table heading will overflow the vertical space allotted for the table heading. The fraction of the page allocated for the table heading may be changed with \tableheadfrac. The num argument to \tableheadfrac should be the decimal fraction of the page used for heading information. The default value is 0.1, meaning that 10% of the page height is reserved for the table heading. It should rarely be necessary to change this value.

2.14.3. Content of the deluxetable

After the table title and column headings are specified, data rows can be entered. Data rows are delimited with the \startdata and \enddata commands. The end of each row is indicated with the standard L^{A}TEX \\ command. Data cells within a row of the table are separated with & (ampersand) characters.

     \startdata

data line\\ cell&cell&cell\\ more data lines

     \enddata

Column alignment within the data columns can be adjusted with the TEX \phantom{string} command, where string can be any character, e.g., \phantom{$\arcmin$}. A blank character of width string is then inserted in the table. Four commands have been predefined for this purpose.

\phn  phantom numeral 0-9
\phd  phantom decimal point
\phs  phantom ± sign
\phm{string} generic phantom

Extra vertical space can be inserted between rows with an optional argument to the \\ command; the argument is a dimension and may be specified in any units that are legitimate in L^{A}TEX.

\\[dimen]

In a table, it may happen that several rows of data are associated with a single object or item; such logical groupings should not be broken across pages. In such cases, the *-form of the \\ command can be used to keep the current line with the next line (“Keep with Next”).

table row\\* next table row

An unlimited number of table rows can be held together in a block in this way.

The journals often require that table cells that contain no data be explicitly marked. This is to differentiate such cells from blank cells, which are frequently interpreted as implicitly repeating the entry in the corresponding cell in the row preceding. Data elements for which there are no data should contain a \nodata command; an appropriate symbol will be placed in that data element.

                                                                 \nodata

Within the deluxetable body, two kinds of “specialty” heads are recognized. A cut-in head is a piece of text that is centered on the table width; it is spaced above and below from the data rows that precede and follow it; there may be rules associated with it, depending on the journal or manuscript style. All of these formatting particulars are managed by the style files. The author need only specify that the text be centered with a \cutinhead command. Similarly, a side head is a piece of text that is left-justified.

                                                                 \cutinhead{text}
                                                                 \sidehead{text}

Table footnotes (more properly, table endnotes) may be used in the deluxetable environment; their use is described in detail in the Section 2.14.5 below.

2.14.4. The table environment

Tables may also be marked and composed using the table environment.

     \begin{table}
     \end{table}

The table environment encloses not only the tabular material but also any title (caption) or footnote information associated with the table.

Titles or captions for tables are indicated with

     \caption{text\label{key}}

Tables will be identified with arabic numerals, e.g., “Table 2”; the identifying text, including the number, is generated automatically by \caption. The author may provide a \label in the caption with a unique key for cross-referencing purposes.

Tabular information is typeset within the tabular environment:

     \begin{tabular}{cols}
     \end{tabular}

where cols specifies the justification for each column. One of the letters ‘l’, ‘c’, or ‘r’ is given for each column, indicating left, center, or right justification. Consult the L^{A}TEX manual (Lamport 1994) for details about using the tabular environment to prepare tables. Each tabular table must appear within a table environment. There should be only one tabular table per table environment. If the journal requests manuscripts with only one table per page, the author may need to insert a \clearpage command after especially short tables.

There is a table line command for use in tabular environments.

     \tableline

This command produces the horizontal rule(s) between the column headings and the body of the table. Authors are discouraged from using \hline. In addition, the use of vertical rules should be avoided.

As with the deluxetable environment, it is possible to override L^{A}TEX’s automatic numbering within the

table environment using \tablenum. \tablenum must

be used inside the table environment.

2.14.5. Table footnotes

AASTEX supports footnotes (endnotes) that are associated with tables; this support applies to both the deluxetable environment and the standard L^{A}TEX table environment. Footnotes for tables are usually identified by lowercase letters rather than numbers. Marking and assigning associated text is achieved

with the \tablenotemark and \tablenotetext com-

mands, which function like the \altaffilmark and \altaffiltext (see Section 2.3 above). The \tablenotetext must be specified after the \end{tabular} or \enddata and before the enclosing \end{table} or

\end{deluxetable}.

                                                                 \tablenotemark{alpha key}
                                                                 \tablenotetext{alpha key}{text}

Note that the alpha key in the \tablenotemark should be the same as the alpha key for the corresponding text. It is the responsibility of the author to make the correspondence correct.

Sometimes authors tabulate materials which have corresponding references, and it may be desirable to associate these references with the table rather than (or in addition to) the formal reference list. Also, authors may occasionally wish to append a short paragraph of explanatory notes that pertain to the entire table, but which are different than the caption. AASTEX offers two special kinds of endnotes for these purposes.

\tablerefs{reference list}

                                                                 \tablecomments{text}

Like the \tablenotetext command, these commands should be specified after the \end{tabular} or \enddata and before the enclosing \end{table} or

                                                      \end{deluxetable}.

2.15. Miscellaneous

2.15.1. Ionic species and chemical bonds

When discussing atomic species, ionization levels can be indicated with the following command.

\ion{element}{level}

The ionization state is specified as the second argument, and should be given as a numeral. For example, Ca III is specified by typing \ion{Ca}{3}.

For single, double, and triple chemical bonds, use the following macros.

                                                                 \sbond
                                                                 \dbond
                                                                 \tbond

2.15.2. Fractions

AASTEX contains commands that permit authors to specify alternate forms for fractions. L^{A}TEX will set fractions in displayed math as built-up fractions; however, it is sometimes desirable to use case fractions in displayed equations. In such instances, one should use \case rather than \frac, although authors submitting manuscripts electronically will generally find it unnecessary to use any markup other than the standard L^{A}TEX \frac.

                                                                Built-up \frac\{1\}\{2\}   \frac{1}{2}
                                                                Case  \case\{1\}\{2\}   \frac{1}{2}

Shilled 1/2  1/2

The AASTEX package also contains a collection of assorted macros for symbols and abbreviations specific to an astronomical context. These are commonly useful and also somewhat difficult for authors to produce themselves because fussy kerning is required. See the symbols pages provided as a separate attachment with this manual. Most of these commands can be used in both running text and math. \lesssim and \gtrsim can only be used in math mode.

2.15.3. Celestial objects

Celestial objects should be marked by the author to facilitate post-processing of articles for the astronomical catalog databases.

\objectname[catalog ID]{text}

This command has no effect in printed manuscripts. When your article contains \objectname commands, the publisher can pass along a list of “objects in article” to database personnel, and software can construct queries to those databases in the online editions.

In the future, it may be useful for the author to specify a particular catalog with an identifier. At this time, however, the optional catalog ID argument is ignored, and need not be supplied.

2.15.4. Hypertext constructs

The \anchor command is a general-purpose hypertext link tag, associating text in the manuscript with the specified resource (href ).

\anchor{href }{text}

href is specified as a full URI, including the scheme: designator (http:, ftp:, etc.).

The \url command supports the special case where an author wishes to express a URL in the text.

\url{text}

The \email command is used to identify email addresses anywhere in the manuscript. The argument is the pertinent email address; please do not prepend the

mailto: part.
\email{address}

This command should be used to indicate authors’ email addresses in author lists at the beginning of manuscripts. The AASTEX v4.0 command, \authoremail, is superseded by \email.

2.16. Concluding the file

The last command in the electronic manuscript file should be the

     \end{document}

command, which appears after all the material in the paper. This command directs the formatter to perform assorted termination activities and finish processing.

3. Style options

The default style option is the manuscript style. This style will produce double-spaced pages printed in a single column at the width of the page.

Two preprint styles are also available. The preprint style offers a single-spaced alternative for manuscripts. This style is similar to the manuscript style, but it is single-spaced, and it is possible to use a smaller typeface. Sometimes, it is desirable for preprints to be set in two columns with running heads. The preprint2 style option achieves this. These two style options are discussed further below.

The eqsecnum style file can be used to modify the way equations are numbered. Normally, equations are numbered sequentially through the entire paper, starting over at “(A1)” if there is an appendix. If eqsecnum appears in the documentclass command, equation numbers will be sequential through each section and will be formatted “(sec-eqn),” where “sec” is the current section number and “eqn” is the number of the equation within that section.

     \documentclass[eqsecnum]{aastex}

A flushrt.sty file option is available for authors that prefer to have their margins left and right justified. The preprint2 style is flush right by default, but the manuscript and preprint styles are ragged right by default.

     \documentclass[eqsecnum,flushrt]{aastex}

3.1. Preprint format

A single-column preprint format can be specified with the preprint style option:

     \documentclass[preprint]{aastex}

The size of the typeface used is under author control by way of L^{A}TEX’s nnpt class options (where nn is 10, 11, or 12). Use of 10 point type is not recommended with the preprint style.

Authors may wish to adjust vertical spacing within a preprint, for instance, double-spacing text while singlespacing tables. Authors who want to alternate between single and double spacing in the manuscript may use the following commands.

     \singlespace
     \doublespace

\singlespace sets the vertical spacing to a smaller value, while \doublespace invokes double-spacing.

3.2. Two-column format

The preprint2 style has the principle function of providing two-column formatting.

     \documentclass[preprint2]{aastex}

It is important to remember that text lines are considerably shorter when two of them are typeset side by side on a page. Long equations, wide tables and figures, and the like, may not typeset in this format without some adjustments. The expenditure of great effort to adapt copy and markup for two-column pages is counterproductive. Remember that the main goal of this package is to produce draft (or referee) format pages; it is the responsibility of the editors and publishers to produce publication-format papers for the journals.

The preprint2 style sets the article’s front matter— the title, author, abstract, and keyword material—on a separate page at full text width. The body of the article is set in a two-column page grid, the appendices in a one-column page grid, and the bibliography in a twocolumn page grid. This manual was prepared using the

preprint2 style.

The author may supply L^{A}TEX’s \twocolumn or \onecolumn commands whenever desired. Be aware, however, that using explicit column-switching commands can cause formatting problems.

4. Additional documentation

The preceding explanation of the markup commands in the AASTEX package has merit for defining syntax, but many authors will prefer to examine the sample papers that are included with the style files. The files of interest are described below.

A comprehensive example employing nearly all of the capabilities of the package (in terms of markup as well as formatting) is in sample.tex. This file is annotated with comments that describe the purpose of most of the markup. sample.tex has three tables: two marked up using the deluxetable environment and another table using the L^{A}TEX table environment.

In table.tex, a complex but short example of the deluxetable environment demonstrates some of the techniques that can be used to generate complex column headings and to align variable-width columns. Here the L^{A}TEX \multicolumn command is used to span a heading over several columns. When \multicolumn is used along

with the \cutinhead command, the \tablecolumns

command must be used to specify the number of columns in the table—otherwise the \cutinhead command will not work properly. This table also makes use of the \phn command to better align some of the columns.

This user guide (aasguide.tex) is also marked up with the AASTEX package, although it is not exemplary as a scientific paper.

Many of the markup commands described in the preceding sections are standard L^{A}TEX commands. The reader who is unfamiliar with their syntax is referred to the L^{A}TEX manual (Lamport 1994), or any of several other guides by Kopka and Daly (1999), Goossens, Mittelbach, and Samarin (1994), and Hahn (1993).

Authors who wish to know the ins and outs of TEX itself should read the TEXbook (Knuth 1984), probably more than once. There is a good deal of information about typography in general in this resource. Many details of mathematical typography are discussed in Mathematics into Type by Swanson (1979).

5. Acknowledgments

AASTEX was designed and written by Chris Biemesderfer in 1988. Substantial revisions were made by Lee Brotzman and Pierre Landau when the package was updated to v4.0. AASTEX was rewritten as a L^{A}TEX 2ε class by Arthur Ogawa when v5.0 was created. The documentation has benefited from revisions by Jeannette Barnes and Sara Zimmerman.

REFERENCES

Abt, H. 1990, ApJ, 357, 1 (editorial)

Adobe Systems, Inc. 1999, PostScript Language Reference Manual (Reading, MA: Addison-Wesley)

Daly, P. 1998, Natural Sciences Citations and References (natbib package documentation)

Goossens, M., Mittelbach, F., and Samarin, A. 1994, The L^{A}TEX Companion (Reading, MA: Addison-Wesley)

Hahn, J. 1993, L^{A}TEX for Everyone (Englewood Cliffs, NJ: Prentice-Hall)

Knuth, D. 1984, The TEXbook (Reading, MA: Addison-Wesley). Revised to cover TEX3, 1991.

Kopka, H. and Daly, P. 1999, A Guide to L^{A}TEX, 3rd edition, (Reading, MA: Addison-Wesley)

Lamport, L. 1994, L^{A}TEX: A Document Preparation System, 2nd edition, (Reading, MA: Addison-Wesley)

Swanson, E. 1979, Mathematics into Type (Providence, RI: American Mathematical Society)

This 2-column preprint was prepared with the AAS L^{A}TEX macros v5.0.